# Dimitrios Kostas [^1] Mathematics subject classification:65M60,65M12,35K55

### v1

@@ Abstract
This study investigates the complex Landau equation, a reaction diffusion system with applications in nonlinear optics and fluid dynamics. The equation’s nonlinear imaginary component introduces rich dynamics and significant computational challenges. We address these challenges using Discontinuous Galerkin (DG) finite element methods. A rigorous stability analysis and a comparative study are performed on three distinct DG schemes: Symmetric Interior Penalty Galerkin (SIPG), Non-symmetric Interior Penalty Galerkin (NIPG), and Incomplete Interior Penalty Galerkin (IIPG). These methods are compared in terms of their stability and computational efficiency.Our numerical analysis and computational results demonstrate that all three discontinuous Galerkin (DG) schemes are stable. However, the Symmetric Interior Penalty Galerkin (SIPG) scheme proves to be the most robust, as its norm remains bounded even in the presence of nonlinear terms—a property not shared by the others. A comparison between the Incomplete Interior Penalty Galerkin (IIPG) and Non-symmetric Interior Penalty Galerkin (NIPG) schemes shows that IIPG has superior stability properties. For high values of the penalty parameter, all methods exhibit similar stability behavior. Our results highlight the suitability of DG methods for simulating complex nonlinear reaction-diffusion systems and provide a practical framework for selecting the most efficient scheme for a given problem.
11footnotetext: Department of Mathematics, and School of Electrical and Computer Engineering National Technical University of Athens, Zografou Campus, Athens 15784, Greece. jimkosta@central.ntua.gr
@@ 1 Introduction
The complex Landau equation, a fundamental model in Nonlinear dynamics, has received limited attention in the existing literature regarding its nonlinear dynamics, with only a handful of studies addressing this topic.To the best of our knowledge, the application of Discontinuous Galerkin (DG) methods—specifically the Symmetric (SIPG), Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin methods—to the complex Landau equation remains unexplored.This study addresses the cretical gap by introducing and analyzing these three Discontinuous Galerkin formulations for the complex Landau equation. Our work establishes novel stability estimates for the proposed schemes and demonstrates their convergence behavior through detailed numerical and theoretical analysis.
@@
The foundational formulation of the interior penalty (IP) method utilizing discontinuous elements was established by Arnold in his seminal work [3], providing the theoretical genesis for subsequent developments, including the widely adopted SIPG, Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin schemes. The flexibility of DG methods in handling complex domains is further evidenced by its successful application in non-overlapping domain decomposition strategies for the incompressible Navier–Stokes equations [12]. Furthermore, efforts to ensure computational reliability and efficiency have driven the development of a posteriori error estimation techniques for DG discretizations of elliptic problems [16], culminating in robust adaptive frameworks that offer convergence guarantees [17]. In the context of computational efficiency, a simplified DG scheme for diffusion-dominated problems has also been proposed, which emphasizes robustness without compromising accuracy [24]. The robustness and accuracy of discontinuous hp-Finite Element Methods in resolving sharp layers in advection–reaction problems, particularly in high-Péclet number regimes, is thoroughly investigated in Houston et al. [14].
@@ 2.1 The Landau equation
In above equation $a,b\in\Re$ are real parameters that significantly influence the system’s temporial and spatial behavior. The presece of the imaginary component in the diffusion term $(1+ib)|u|^{2}u$ gives rise to nonlinear frequency modulation, in which the wave frequncy becomes ampiltude-dependent.
The complex coefficients play a crucial role in determining the qualitive nature of the solutions. In particular, they leafd to a variety of rich spatiotemporial including
@@
The part b ios often assumed to be non-negative to avoid unphysical behaviour b¿0.
@@
The bilinear form $a_{h}(u,v)$ which defined the discrete operator used in our system 2.2 2.2 is explicitly introduced in the equation 2.13 and reflects the interrior penalty stucture essentially for the Discontinuous Galerkin framework.
@@ 2.3 Existence and Uniqueness of the Continuous and Discrete Problems
The Landau equation, in its regularized form, is a nonlinear parabolic partial differential equation (PDE) of Fokker–Planck type. Under standard assumptions on the regularity and decay of the initial data, and with appropriate smoothing of the collision kernel, classical results guarantee the existence and uniqueness of weak solutions in suitable spaces $L^{2}(0,T,H^{1}(\Omega))$ Reference [1] Discusses the ,derivation, justification, and weak solution of the Landau equation from physical models. A comprehensive of well-posedness and convergence results for discretized (numerical) versions of the Landau equation is provided in [11], while [21] establishes coercivity and regularity results for the linearized Landau operator.
The fully discrete problem, where we approximate the Landau equation using a discontinuous Galerkin (DG) or continuous Galerkin (CG) finite element scheme, we consider the discrete formulation in a finite-dimensional function space. Due to the finite dimensionality and the local Lipschitz continuity of the nonlinear discrete operator, the discrete problem admits a unique solution that is local in time . Furthermore, under standard coercivity and boundedness assumptions on the discrete bilinear forms—particularly for DG schemes equipped with appropriate penalty parameters—energy estimates can be derived to ensure global-in-time existence and uniqueness of the discrete solution. A brief stability analysis further confirms that the numerical method is well-posed and robust under mesh refinement.
@@ 3 Stability Estimate
Step 1 We begin with bounding the bilinear form $a_{h}(u_{h},v_{h})$ which includes volume and penalty interface terms that ensure coercivity in the DG norm.This form has penalty terms.
@@
Let $\Omega\subset\Re^{d}$be a bounded Lipschitz domain, and ${T}^{n}_{i=1}$ be a comforming shape regular partition of $\Omega$ into the disjoint element such that $\Omega=\cup_{i=1}^{N}T^{i}$ with $T_{i}\cap T_{j}=$ for $i\neq j$ and time interval be discretized into a partition in time $0=t^{1}<t^{2}<...<t^{N}=T.$ Assume that the bilinear form $a_{h}(u_{h},v),$ defined in equation 2.13, corresponds to a time discretization using theta scheme ,and a special discretization based on one of the interrior penalty discontinuous Galerkin methods (SIPG),(NIPG),(IIPG) For penalty parameters $\sigma$ large enough If it holds the 2.12 Poincare’ inequality. Then there exists constants $\beta>0,c\geq 0$ and $C>0$ independent of the mesh size h and time step $\Delta t$, that the bilinear term satisfies the following estimates:

### v2

@@ Abstract
This study investigates the complex Landau equation, a reaction diffusion system with applications in nonlinear optics and fluid dynamics. The equation’s nonlinear imaginary component introduces rich dynamics and significant computational challenges. We address these challenges using Discontinuous Galerkin (DG) finite element methods. A rigorous stability analysis and a comparative study are performed on three distinct DG schemes: Symmetric Interior Penalty Galerkin (SIPG), Non-symmetric Interior Penalty Galerkin (NIPG), and Incomplete Interior Penalty Galerkin (IIPG). These methods are compared in terms of their stability and computational efficiency.Our numerical analysis and computational results demonstrate that all three discontinuous Galerkin (DG) schemes are stable. However, the Symmetric Interior Penalty Galerkin (SIPG) scheme proves to be the most robust, as its norm remains bounded even in the presence of nonlinear terms. A property not shared by the others. A comparison between the Incomplete Interior Penalty Galerkin (IIPG) and Non-symmetric Interior Penalty Galerkin (NIPG) schemes shows that IIPG has superior stability properties. For high values of the penalty parameter, all methods exhibit similar stability behavior. Our results highlight the suitability of DG methods for simulating complex nonlinear reaction-diffusion systems and provide a practical framework for selecting the most efficient scheme for a given problem.
11footnotetext: The work completed National Technical University of Athens, Zografou Campus, Athens 15784, Greece. jimkosta@central.ntua.gr Now ind. researcher
@@ 1 Introduction
The complex Landau equation, a fundamental model in Nonlinear dynamics, has received limited attention in the existing literature regarding its nonlinear dynamics, with only a handful of studies addressing this topic.To the best of our knowledge, the application of Discontinuous Galerkin (DG) methods—specifically the Symmetric (SIPG), Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin methods—to the complex Landau equation remains unexplored.This study addresses the critical gap by introducing and analyzing these three Discontinuous Galerkin formulations for the complex Landau equation. Our work establishes novel stability estimates for the proposed schemes and demonstrates their convergence behavior through detailed numerical and theoretical analysis.
@@
The foundational formulation of the interior penalty (IP) method utilizing discontinuous elements was established by Arnold in his seminal work [3], providing the theoretical genesis for subsequent developments, including the widely adopted SIPG, Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin schemes. The flexibility of DG methods in handling complex domains is further evidenced by its successful application in non-overlapping domain decomposition strategies for the incompressible Navier–Stokes equations [12]. Furthermore, efforts to ensure computational reliability and efficiency have driven the development of A-posteriori error estimation techniques for DG discretizations of elliptic problems [16], culminating in robust adaptive frameworks that offer convergence guarantees [17]. In the context of computational efficiency, a simplified DG scheme for diffusion-dominated problems has also been proposed, which emphasizes robustness without compromising accuracy [24]. The robustness and accuracy of discontinuous hp-Finite Element Methods in resolving sharp layers in reaction diffusion problems, particularly in high-Péclet number regimes, is thoroughly investigated in Houston et al. [14].
@@ 2.1 The Landau equation
In above equation $a,b\in\Re$ are real parameters that significantly influence the system’s temporal and spatial behavior. The presece of the imaginary component in the diffusion term $(1+ib)|u|^{2}u$ gives rise to nonlinear frequency modulation, in which the wave frequency becomes ampiltude-dependent.
The complex coefficients play a crucial role in determining the qualitive nature of the solutions. In particular, they lead to a variety of rich spatiotemporial including
@@
The part b ios often assumed to be non-negative to avoid unphysical behavior b¿0.
@@
The bilinear form $a_{h}(u,v)$ which defined the discrete operator used in our system 2.2 2.2 is explicitly introduced in the equation 2.13 and reflects the interior penalty stucture essentially for the Discontinuous Galerkin framework.
@@ 2.3 Existence and Uniqueness of the Continuous and Discrete Problems
The Landau equation, in its regularized form, is a nonlinear parabolic partial differential equation (PDE) of Fokker–Planck type. Under standard assumptions on the regularity and decay of the initial data, and with appropriate smoothing of the collision kernel, classical results guarantee the existence and uniqueness of weak solutions in suitable spaces $L^{2}(0,T,H^{1}(\Omega))$ Reference [1] Discusses the ,derivation, justification, and weak solution of the Landau equation from physical models. A comprehensive of well-possessiveness and convergence results for discretized (numerical) versions of the Landau equation is provided in [11], while [21] establishes results for the linearized Landau operator.
The fully discrete problem, where we approximate the Landau equation using a discontinuous Galerkin (DG) or continuous Galerkin (CG) finite element scheme, we consider the discrete formulation in a finite-dimensional function space. Due to the finite dimensionality and the local Lipschitz continuity of the nonlinear discrete operator, the discrete problem admits a unique solution that is local in time . Furthermore, under standard coercivity and boundedness assumptions on the discrete bi-linear forms—particularly for DG schemes equipped with appropriate penalty parameters—energy estimates can be derived to ensure global-in-time existence and uniqueness of the discrete solution. A brief stability analysis further confirms that the numerical method is well-posed and robust under mesh refinement.
@@ 3 Stability Estimate
Step 1 We begin with bounding the bi-linear form $a_{h}(u_{h},v_{h})$ which includes volume and penalty interface terms that ensure coercivity in the DG norm.This form has penalty terms.
@@
Let $\Omega\subset\Re^{d}$be a bounded Lipschitz domain, and ${T}^{n}_{i=1}$ be a comforming shape regular partition of $\Omega$ into the disjoint element such that $\Omega=\cup_{i=1}^{N}T^{i}$ with $T_{i}\cap T_{j}=$ for $i\neq j$ and time interval be discretized into a partition in time $0=t^{1}<t^{2}<...<t^{N}=T.$ Assume that the bilinear form $a_{h}(u_{h},v),$ defined in equation 2.13, corresponds to a time discretization using theta scheme ,and a special discretization based on one of the interrior penalty discontinuous Galerkin methods (SIPG),(NIPG),(IIPG) For penalty parameters $\sigma$ large enough If it holds the 2.12 Poincare’ inequality. Then there exists constants $\beta>0,c\geq 0$ and $C>0$ independent of the mesh size h and time step $\Delta t$, that the bi-linear term satisfies the following estimates: